OpenStudy (anonymous):

a first number is 4 less than a second number. Four times the first number is 6 more than twice the second. find the number

OpenStudy (anonymous):

f=s-4 4f=2s+6 Does this help?

OpenStudy (anonymous):

Hint: s=11

OpenStudy (anonymous):

In case you didn't get it, f=first number s=second number

OpenStudy (owlcoffee):

replace f=s-4 on 4f=2s+6. You'll end up with: 4(s-4)=2s+6 just solve it for s :)